# An In-depth Evaluation of GPT-4 in Sentence Simplification with Error-based Human Assessment

### v1

@@ Abstract.
Sentence simplification, which rewrites a sentence to be easier to read and understand, is a promising technique to help people with various reading difficulties. With the rise of advanced large language models (LLMs), evaluating their performance in sentence simplification has become imperative. Recent studies have used both automatic metrics and human evaluations to assess the simplification abilities of LLMs. However, the suitability of existing evaluation methodologies for LLMs remains in question. First, the suitability of current automatic metrics on LLMs’ simplification evaluation is still uncertain. Second, current human evaluation approaches in sentence simplification often fall into two extremes: they are either too superficial, failing to offer a clear understanding of the models’ performance, or overly detailed, making the annotation process complex and prone to inconsistency, which in turn affects the evaluation’s reliability. To address these problems, this study provides in-depth insights into LLMs’ performance while ensuring the reliability of the evaluation. We design an error-based human annotation framework to assess the GPT-4’s simplification capabilities. Results show that GPT-4 generally generates fewer erroneous simplification outputs compared to the current state-of-the-art. However, LLMs have their limitations, as seen in GPT-4’s struggles with lexical paraphrasing. Furthermore, we conduct meta-evaluations on widely used automatic metrics using our human annotations. We find that while these metrics are effective for significant quality differences, they lack sufficient sensitivity to assess the overall high-quality simplification by GPT-4.
large language models, GPT-4, evaluation, sentence simplification
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Computing methodologies Natural language generation
@@ 1. Introduction
Sentence simplification automatically rewrites sentences to make them easier to read and understand by modifying their wording and structures, without changing their meanings. It helps people with reading difficulties, such as non-native speakers (Paetzold, 2016), individuals with aphasia (Carroll et al., 1999), dyslexia (Rello et al., 2013b, a), or autism (Barbu et al., 2015). Previous studies often employed a sequence-to-sequence model, then has been enhanced by integrating various sub-modules into it (Zhang and Lapata, 2017; Zhao et al., 2018; Nishihara et al., 2019; Martin et al., 2020). Recent developments have seen the rise of large language models (LLMs). Notably, the ChatGPT families released by OpenAI demonstrate exceptional general and task-specific abilities  (OpenAI, 2023; Wang et al., 2023; Liu et al., 2023), and sentence simplification is not an exception.
Some studies (Feng et al., 2023; Kew et al., 2023) have begun to evaluate LLMs’ performance in sentence simplification, including both automatic scoring and conventional human evaluations where annotators assess the levels of fluency, meaning preservation, and simplicity (Kriz et al., 2019; Jiang et al., 2020; Alva-Manchego et al., 2021; Maddela et al., 2021), or identify common edit operations (Alva-Manchego et al., 2017). However, these studies face limitations and challenges. Firstly, it is unclear whether the current automatic metrics are suitable for evaluating simplification abilities of LLMs. Although these metrics have demonstrated variable effectiveness across conventional systems (e.g., rule-based (Sulem et al., 2018b), statistical machine translation-based (Wubben et al., 2012; Xu et al., 2016), and sequence-to-sequence model-based simplification (Zhang and Lapata, 2017; Martin et al., 2020)) through their correlation with human evaluations (Alva-Manchego et al., 2021), their suitability for LLMs has yet to be explored, thereby their effectiveness in assessing LLMs’ simplifications are uncertain. Secondly, given the general high performance of LLMs, conventional human evaluations may be too superficial to capture the subtle yet critical aspects of simplification quality. This lack of depth undermines the interpretability when evaluating on LLMs. Recently, Heineman et al. (Heineman et al., 2023) proposed a detailed human evaluation framework for LLMs, categorizing $21$ linguistically based success and failure types. However, their linguistics-based approach appears to be excessively intricate and complex, resulted in low consistency among annotators, thus raising concerns about the reliability of the evaluation. The trade-off between interpretability and reliability underscores the necessity for a more balanced approach.
@@
We applied our error-based human evaluation framework to evaluate the performance of GPT-4111We used the ‘gpt-4-0613’ and accessed it via OpenAI’s APIs. in English sentence simplification, using prompt engineering on three representative datasets on sentence simplification: Turk (Xu et al., 2016), ASSET (Alva-Manchego et al., 2020), and Newsela (Xu et al., 2015). The results indicate that GPT-4 generally surpasses the previous state-of-the-art (SOTA) in performance; GPT-4 tends to generate fewer erroneous simplification outputs and better preserve the original meaning, while maintaining comparable levels of fluency and simplicity. However, LLMs have their limitations, as seen in GPT-4’s struggles with lexical paraphrasing. The meta-evaluation results show that existing automatic metrics face difficulties in assessing simplification generated by GPT-4. That is, these metrics are able to differentiate significant quality differences but not sensitive enough for overall high-quality simplification outputs.
@@ 2. Related Work
Our study evaluates the performance of the advanced LLM, namely GPT-4, in sentence simplification by comparing it against the SOTA supervised simplification model. This section includes a review of current evaluations of LLMs in this domain, along with an overview of the SOTA supervised simplification models.
@@ 2.1. Evaluation of LLM-based Simplification
In sentence simplification, some studies attempted to assess the performance of LLMs. For example, Feng et al. (Feng et al., 2023) evaluated the performance of prompting ChatGPT and GPT-3.5, later Kew et al. (Kew et al., 2023) compared $44$ LLMs varying in size, architecture, pre-training methods, and with or without instruction tuning. Additionally, Heineman et al. (Heineman et al., 2023) proposed a detailed human evaluation framework for LLMs, categorizing 21 linguistically based success and failure types. Their findings indicate that OpenAI’s LLMs generally surpass the previous SOTA supervised simplification models.
However, these studies have three primary limitations. First, there has not been a comprehensive exploration into the capabilities of the most advanced ChatGPT model to date, i.e., GPT-4. Second, these studies do not adequately explore prompt variation, employing uniform prompts with few-shot examples across datasets without considering their unique features in simplification strategies. This may underutilize the potential of LLMs, which are known to be prompt-sensitive. Third, the human evaluations conducted are inadequate. Such evaluations are crucial, as automatic metrics often have blind spots and may not always be entirely reliable (He et al., 2023). Human evaluations in these studies often rely on shallow ratings or edit operation identifications to evaluate a narrow range of simplification outputs. These methods risk being superficial, overlooking intricate features. In contrast, Heineman et al.’s linguistics-based approach (Heineman et al., 2023) appears to be excessively intricate and complex, resulting in low consistency among annotators, thus raising concerns about the reliability of the evaluations. Our study aims to bridge these gaps, significantly enhancing the utility of ChatGPT models through comprehensive prompt engineering processes, and incorporating elaborate human evaluations while ensuring reliability.
@@ 2.2. SOTA Supervised Simplification Models
Traditional NLP methods heavily relied on task-specific models, which involve adapting pre-trained language models for various downstream applications. In sentence simplification, Martin et al. (Martin et al., 2022) introduced the MUSS model by fine-tuning BART (Lewis et al., 2020) with labeled sentence simplification datasets and/or mined paraphrases. Similarly, Sheang et al. (Sheang and Saggion, 2021) fine-tuned T5 (Raffel et al., 2020), which is called Control-T5 in this study, achieving SOTA performance on two representative datasets: Turk and ASSET. These models leverage control tokens, which have been initially introduced by ACCESS (Martin et al., 2020), to modulate attributes like length, lexical complexity, and syntactic complexity during simplification. This approach allows any sequence-to-sequence model to adjust these attributes by conditioning on simplification-specific tokens, facilitating strategies that aim to shorten sentences or reduce their lexical and syntactic complexity. Our study employs Control-T5 as a SOTA model and compares it to GPT-4 in sentence simplification.
@@ 3. Datasets
In this study, we employed standard datasets for English sentence simplification, as detailed below. For replicating the SOTA supervised model, namely, Control-T5, we used the same training datasets as the original paper. Meanwhile, the evaluation datasets were used to assess the performance of our models.
@@ 3.1. Training Datasets
We utilized training sets from two datasets: WikiLarge (Zhang and Lapata, 2017) and Newsela (Xu et al., 2015; Zhang and Lapata, 2017). WikiLarge consists of $296k$ complex-simple sentence pairs automatically extracted from English Wikipedia and Simple English Wikipedia by sentence alignment. Introduced by Xu et al. (Xu et al., 2015), Newsela originates from a collection of news articles accompanied by simplified versions written by professional editors. It was subsequently aligned from article-level to sentence-level, resulting in approximately $94k$ complex-simple sentence pairs. In our study, we utilize the training split of the Newsela dataset made by Zhang and Lapata (Zhang and Lapata, 2017).
@@ 3.2. Evaluation Datasets
We used validation and test sets from the three standard datasets on English sentence simplification. [^2] These validation sets were used for prompt engineering on GPT-4. Table 1 shows the numbers of complex-simple sentence pairs in these sets. These datasets have distinctive features due to differences in simplification strategies as summarized below.
@@
  • ASSET (Alva-Manchego et al., 2020): This dataset uses the same $2,359$ source sentences as the Turk dataset. It differs from Turk by aiming at rewriting sentences with more diverse transformations, i.e., paraphrasing, deleting phrases, and splitting a sentence, and provides $10$ simplified references written by crowd-workers.
@@
  • Newsela (Xu et al., 2015; Zhang and Lapata, 2017): This is the same Newsela dataset described in Section 3.1. We utilized its validation and test splits, totaling $2,206$ sentence pairs. After careful observation, we found that deletions of words, phrases, and clauses predominantly characterize the Newsela dataset.
@@
Table 1. Number of complex-simple sentence pairs in the validation and test sets of each dataset
@@
4. Models
To enhance the performance of GPT-4 in sentence simplification, we undertook extensive prompt engineering effort. We also replicated the SOTA supervised model, Control-T5, for comparative analysis with GPT-4. Throughout our optimization efforts, we employed SARI (Xu et al., 2016), which is a widely recognized statistic-based metric for evaluating sentence simplification. [^3] Our meta-evaluation in Section 7 confirms that SARI score aligns with human evaluation. SARI evaluates a simplification model by comparing its outputs against references and source sentences, focusing on the words that are added, kept, and deleted. Its values range from 0 to $100$, with higher values indicating better performance.
4.1. GPT-4 with Prompt Engineering
Scaling pre-trained language models, such as increasing model and data size, LLMs enhance their capacity for downstream tasks. Unlike earlier models that required fine-tuning, these LLMs can be effectively prompted with zero- or few-shot examples for task-solving.
4.1.1. Design
Aiming to optimize GPT-4’s sentence simplification capabilities, we conducted prompt engineering based on three principal components:
  • Varied Number of References: We experimented with single or multiple (namely, three) simplification references used in the examples. For Turk and ASSET, which are multi-reference datasets, we manually selected one high-quality reference from their multiple references. Newsela, which is basically a single-reference dataset, offers multiple simplification levels for the same source sentences. For this dataset, we extracted references targeting different simplicity levels of the same source sentence as multiple references.
We integrated these components into prompts, resulting in the creation of $15$ variations (Figure 1). These prompts were then applied to each validation set, excluding selected examples. Prompts that achieved the highest SARI scores were designated as ‘Best Prompts’, which are summarized in Table 2. For more detailed information, refer to the Appendix A.1. Following this, we used the best prompts to generate simplification outputs from the respective test sets.
4.1.2. Effect of Prompt Engineering
Prompt engineering demonstrates its effectiveness. As shown in Table 2, across three validation sets, prompts with the highest SARI scores significantly outperformed those with the lowest, achieving scores of $8.3$ for Turk, $4.5$ for ASSET, and $3.6$ for Newsela. Moreover, results reveal a direct alignment between the best prompt’s instructional style and its respective dataset. These top-performing prompts all use a few-shot examples of three. The optimal number of simplification references varies; Turk and ASSET show strong results with a single reference, whereas Newsela benefits from multiple references, likely due to the intricacies involved in ensuring meaning is preserved amidst deletions. Overall, prompt engineering notably enhances GPT-4’s sentence simplification output, as evidenced by the significant increase in SARI scores.
Figure 1. GPT-4 prompt engineering; the identified best prompt on the ASSET validation set is illustrated here.
Table 2. Prompt engineering impact on SARI scores
@@ 4.1.2. Effect of Prompt Engineering
Turk $8.3$ Turk style + Few-shot + Single ref
ASSET $4.5$ ASSET style + Few-shot + Single ref
Newsela $3.6$ Newsela style + Few-shot + Multi refs

### v2

# An In-depth Evaluation of Large Language Models in Sentence Simplification with Error-based Human Assessment

@@ Abstract.
Sentence simplification, which rewrites a sentence to be easier to read and understand, is a promising technique to help people with various reading difficulties. With the rise of advanced large language models (LLMs), evaluating their performance in sentence simplification has become imperative. Recent studies have used both automatic metrics and human evaluations to assess the simplification abilities of LLMs. However, the suitability of existing evaluation methodologies for LLMs remains in question. First, the suitability of current automatic metrics on LLMs’ simplification evaluation is still uncertain. Second, current human evaluation approaches in sentence simplification often fall into two extremes: they are either too superficial, failing to offer a clear understanding of the models’ performance, or overly detailed, making the annotation process complex and prone to inconsistency, which in turn affects the evaluation’s reliability. To address these problems, this study provides in-depth insights into LLMs’ performance while ensuring the reliability of the evaluation. We design an error-based human annotation framework to assess the LLMs’ simplification capabilities. We select both closed-source and open-source LLMs, including GPT-4, Qwen2.5-72B, and Llama-3.2-3B. We believe that these models offer a representative selection across large, medium, and small sizes of LLMs. Results show that GPT-4 generally generates fewer erroneous simplification outputs compared to the current state-of-the-art. However, LLMs have their limitations, as seen in GPT-4’s struggles with lexical paraphrasing. Results show that LLMs generally generate fewer erroneous simplification outputs compared to the previous state-of-the-art. However, LLMs have their limitations, as seen in GPT-4’s and Qwen2.5-72B’s struggle with lexical paraphrasing. Furthermore, we conduct meta-evaluations on widely used automatic metrics using our human annotations. We find that these metrics lack sufficient sensitivity to assess the overall high-quality simplifications, particularly those generated by high-performance LLMs.
large language models, evaluation, sentence simplification
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Computing methodologies Natural language generation
@@ 1. Introduction
Sentence simplification automatically rewrites sentences to make them easier to read and understand by modifying their wording and structures, without changing their meanings. It helps people with reading difficulties, such as non-native speakers (Paetzold, 2016), individuals with aphasia (Carroll et al., 1999), dyslexia (Rello et al., 2013b, a), or autism (Barbu et al., 2015). Previous studies often employed a sequence-to-sequence model, which was then enhanced by integrating various sub-modules into it (Zhang and Lapata, 2017; Zhao et al., 2018; Nishihara et al., 2019; Martin et al., 2020). Recent developments have seen the rise of large language models (LLMs). Among them, the closed-source ChatGPT families released by OpenAI demonstrate exceptional general and task-specific abilities (OpenAI, 2023; Wang et al., 2023a; Liu et al., 2023), and sentence simplification is not an exception. On the open-source front, LLMs such as the Llama family by Meta (et al, 2023a) and the Qwen family by Alibaba Cloud (et al, 2023b) stand out as prominent representatives, showing competitive performance.
Some studies (Feng et al., 2023; Kew et al., 2023) have begun to evaluate LLMs’ performance in sentence simplification, including both automatic scoring and conventional human evaluations where annotators assess the levels of fluency, meaning preservation, and simplicity (Kriz et al., 2019; Jiang et al., 2020; Alva-Manchego et al., 2021; Maddela et al., 2021), or identify common edit operations (Alva-Manchego et al., 2017). However, these studies face limitations and challenges. Firstly, it is unclear whether the current automatic metrics are suitable for evaluating the simplification abilities of LLMs. Although these metrics have demonstrated variable effectiveness across conventional systems (e.g., rule-based (Sulem et al., 2018b), statistical machine translation-based (Wubben et al., 2012; Xu et al., 2016), and sequence-to-sequence model-based simplification (Zhang and Lapata, 2017; Martin et al., 2020)) through their correlation with human evaluations (Alva-Manchego et al., 2021), their suitability for LLMs has yet to be explored, thereby their effectiveness in assessing LLMs’ simplifications are uncertain. Secondly, given the general high performance of LLMs, conventional human evaluations may be too superficial to capture the subtle yet critical aspects of simplification quality. This lack of depth undermines the interpretability when evaluating LLMs. Recently, Heineman et al. (Heineman et al., 2023) proposed a detailed human evaluation framework for LLMs, categorizing $21$ linguistically based success and failure types. However, their linguistics-based approach appears to be excessively intricate and complex, resulting in low consistency among annotators, thus raising concerns about the reliability of the evaluation. The trade-off between interpretability and reliability underscores the necessity for a more balanced approach.
@@
We apply our error-based human evaluation framework to evaluate the performance of GPT-4111We used the ‘gpt-4-0613’ and accessed it via OpenAI’s APIs., Qwen2.5-72B, and Llama-3.2-3B [^2] We used the ‘Qwen2.5-72B-Instruct’ and ‘Llama-3.2-3B-Instruct’. We ran the two models using Transformers library(et al, 2019). in English sentence simplification. We believe that these models offer a representative selection across large, medium, and small sizes of LLMs. We use prompt engineering and evaluate models on four representative datasets on sentence simplification: Turk (Xu et al., 2016), ASSET (Alva-Manchego et al., 2020a), Newsela (Xu et al., 2015), and SimPA (Scarton et al., 2018). Figure 1 illustrates the overview of our evaluation pipeline. Our key findings are summarized as follows:
  • LLMs generally surpass the previous state-of-the-art (SOTA) in performance; LLMs tend to generate fewer erroneous simplification outputs and better preserve the original meaning, while maintaining comparable levels of fluency and simplicity.
  • Among the LLMs, GPT-4 and Qwen2.5-72B surpass Llama-3.2-3B, with Qwen2.5-72B generating fewer errors than GPT-4. This implies the strong potential of medium-sized LLMs in simplification tasks.
  • However, larger LLMs have their limitations, as seen in GPT-4 and Qwen2.5-72B’s struggles with lexical paraphrasing.
  • The meta-evaluation reveals that existing automatic metrics struggle to effectively differentiate between high- and low-quality simplifications labeled by human, particularly when evaluating the overall high-quality outputs of GPT-4 and Qwen2.5-72B.
Figure 1. Overview of our methodology: sequential evaluation pipeline with human assessment example
@@ 2. Related Work
Our study evaluates the performance of representative closed-source and open-source LLMs of varying sizes in sentence simplification by comparing them against the SOTA supervised simplification model. This section includes a review of current evaluations of LLMs in this domain, along with an overview of the SOTA supervised simplification model.
@@ 2.1. Evaluation of LLM-based Simplification
In sentence simplification, some studies attempted to assess the performance of LLMs. For example, Feng et al. (Feng et al., 2023) evaluated the performance of prompting ChatGPT and GPT-3.5; later, Kew et al. (Kew et al., 2023) compared $44$ LLMs varying in size, architecture, pre-training methods, and with or without instruction tuning. Additionally, Heineman et al. (Heineman et al., 2023) proposed a detailed human evaluation framework for LLMs, categorizing 21 linguistically based success and failure types. Their findings indicate that OpenAI’s LLMs generally surpass the previous SOTA supervised simplification models.
However, these studies have three primary limitations. First, there has not been a comprehensive exploration into the capabilities of the most advanced closed-source and open-source models to date, i.e., GPT-4 and Llama-3. Second, these studies do not adequately explore prompt variation, employing uniform prompts with few-shot examples across datasets without considering their unique features in simplification strategies. This may underutilize the potential of LLMs, which are known to be prompt-sensitive. Third, the human evaluations conducted are inadequate. Such evaluations are crucial, as automatic metrics often have blind spots and may not always be entirely reliable (He et al., 2023). Human evaluations in these studies often rely on shallow ratings or edit operation identifications to evaluate a narrow range of simplification outputs. These methods risk being superficial, overlooking intricate features. In contrast, Heineman et al.’s linguistics-based approach (Heineman et al., 2023) appears to be excessively intricate and complex, resulting in low consistency among annotators, thus raising concerns about the reliability of the evaluations. Our study aims to bridge these gaps, significantly enhancing the utility of LLMs through comprehensive prompt engineering processes, and incorporating elaborate human evaluations while ensuring reliability.
@@ 2.2. SOTA Supervised Simplification Models
Traditional NLP methods heavily relied on task-specific models, which involve adapting pre-trained language models for various downstream applications. In sentence simplification, Martin et al. (Martin et al., 2022) introduced the MUSS model by fine-tuning BART (Lewis et al., 2020) with labeled sentence simplification datasets and/or mined paraphrases. Similarly, Sheang et al. (Sheang and Saggion, 2021) fine-tuned T5 (Raffel et al., 2020), which is called Control-T5 in this study, achieving SOTA performance on two representative datasets: Turk and ASSET. These models leverage control tokens, which were initially introduced by ACCESS (Martin et al., 2020), to modulate attributes like length, lexical complexity, and syntactic complexity during simplification. This approach allows any sequence-to-sequence model to adjust these attributes by conditioning on simplification-specific tokens, facilitating strategies that aim to shorten sentences or reduce their lexical and syntactic complexity. Our study employs Control-T5 as the previous SOTA model and compares it to LLMs in sentence simplification.
@@ 3. Datasets
In this study, we employ standard datasets for English sentence simplification, as detailed below. For replicating the SOTA supervised model, namely, Control-T5, we use the same training datasets as the original paper. Meanwhile, the evaluation datasets are used to assess the performance of our models.
@@ 3.1. Training Datasets
We use training sets from two datasets: WikiLarge (Zhang and Lapata, 2017) and Newsela (Xu et al., 2015; Zhang and Lapata, 2017). WikiLarge consists of $296k$ complex-simple sentence pairs automatically extracted from English Wikipedia and Simple English Wikipedia by sentence alignment. Introduced by Xu et al. (Xu et al., 2015), Newsela originates from a collection of news articles accompanied by simplified versions written by professional editors. It was subsequently aligned from article-level to sentence-level, resulting in approximately $94k$ complex-simple sentence pairs. In our study, we utilize the training split of the Newsela dataset made by Zhang and Lapata (Zhang and Lapata, 2017).
@@ 3.2. Evaluation Datasets
We use validation and test sets from four datasets on English sentence simplification. [^3] Validation sets from Turk, ASSET, and Newsela were used for prompt engineering on GPT-4. Table 2 shows the numbers of complex-simple sentence pairs in these sets. These datasets have distinctive features due to differences in simplification strategies and as summarized below.
@@
  • ASSET (Alva-Manchego et al., 2020a): This dataset uses the same $2,359$ source sentences as the Turk dataset. It differs from Turk by aiming at rewriting sentences with more diverse transformations, i.e., paraphrasing, deleting phrases, and splitting a sentence, and provides $10$ simplified references written by crowd-workers.
@@
  • Newsela (Xu et al., 2015; Zhang and Lapata, 2017): This is the same Newsela dataset described in Section 3.1. We utilize its validation and test splits, totaling $2,206$ sentence pairs. After careful observation, we found that deletions of words, phrases, and clauses predominantly characterize the Newsela dataset.
  • SimPA (Scarton et al., 2018): This dataset originated from the public administration domain. It contains $1,100$ original sentences with two versions of simplified sentences: (1) lexical simplifications (2) lexical and syntactic simplifications. We select the second version for its diverse transformations. Note that SimPA does not provide validation/test splits. We use this dataset exclusively as a test set, excluding three sentence pairs reserved for 3-shot examples.
4. Models
To enhance the performance of LLMs in sentence simplification, we undertook prompt engineering on GPT-4 across validation datasets, and adapted the prompts for GPT-4, Qwen2.5-72B, and Llama-3.2-3B models. For SimPA, which shares ASSET’s diverse transformation characteristic, we reused the optimized instruction for ASSET without additional prompt engineering. We also replicated the SOTA supervised model, Control-T5, for comparative analysis with LLMs. Throughout our optimization efforts, we employed SARI (Xu et al., 2016), which is a widely recognized statistic-based metric for evaluating sentence simplification. [^4] Our meta-evaluation in Section 7 confirms that SARI score aligns with human evaluation. SARI evaluates a simplification model by comparing its outputs against references and source sentences, focusing on the words that are added, kept, and deleted. Its values range from 0 to $100$, with higher values indicating better performance.
4.1. LLMs with Prompt Engineering
By scaling pre-trained language models, such as increasing model and data size, LLMs enhance their capacity for downstream tasks. Unlike earlier models that required fine-tuning, these LLMs can be effectively prompted with zero- or few-shot examples for task-solving. Previous research has looked into different prompting techniques for various tasks. For example, prompt chaining has been studied for summarization (Sun et al., 2024). In the area of named entity recognition, the use of special tokens (like @@##) has been found to enhance entity identification (Wang et al., 2023b). In our study, we utilize existing human annotation guidelines from these datasets. This method has been employed in earlier studies, demonstrating its effectiveness (Mishra et al., 2022; Sainz et al., 2024).
4.1.1. Design
Aiming to optimize LLMs’ sentence simplification capabilities, we conducted prompt engineering on GPT-4 based on three principal components:
  • Varied Number of References: We experimented with a single or multiple (namely, three) simplification references used in the examples. For Turk and ASSET, which are multi-reference datasets, we manually selected one high-quality reference from their multiple references. Newsela, which is basically a single-reference dataset, offers multiple simplification levels for the same source sentences. For this dataset, we extracted references targeting different simplicity levels of the same source sentence as multiple references.
We integrated these components into prompts, resulting in the creation of $15$ variations. These prompts were then applied to each validation set, excluding selected examples. Prompts that achieved the highest SARI scores were designated as ‘Best Prompts’, which are summarized in Table 2. For more detailed information, refer to the Appendix A.1. Following this, we used the best prompts to generate simplification outputs from the respective test sets.
4.1.2. Effect of Prompt Engineering
Prompt engineering demonstrates its effectiveness. As shown in Table 2, across three validation sets, prompts with the highest SARI scores significantly outperform those with the lowest, achieving scores of $8.3$ for Turk, $4.5$ for ASSET, and $3.6$ for Newsela. Moreover, results reveal a direct alignment between the best prompt’s instructional style and its respective dataset. These top-performing prompts all use a few-shot examples of three. The optimal number of simplification references varies; Turk and ASSET show strong results with a single reference, whereas Newsela benefits from multiple references, likely due to the intricacies involved in ensuring that meaning is preserved amidst deletions. Again, SimPA was not included in the prompt engineering process. Instead, we directly applied the instruction from ASSET, accompanied by 3-shot examples with single references from SimPA itself, given the similarity between the two datasets in their emphasis on diverse transformations. Overall, prompt engineering notably enhances GPT-4’s sentence simplification output, as evidenced by the significant increase in SARI.
@@ 4.1.2. Effect of Prompt Engineering
Table 1. Number of complex-simple sentence pairs in the validation and test sets of each dataset.
@@
SimPA 00 $1,097$
Table 2. The Impact of Prompt Engineering on SARI Scores: Few-Shot (FS), Single Reference (SR), and Multi-Reference (MR)
@@
Turk $8.3$ Turk style + FS + SR
ASSET $4.5$ ASSET style + FS + SR
Newsela $3.6$ Newsela style + FS + MR